A slef-propelled vehicle of mass m , whose  engine dilivers a constant power  P has an accerlation a=(P/mv) .(assume that there is  no friction) In  order to increase its velocity from  V1 to V2, the distance it has to travel is will be
 

Grade:12th Pass

1 Answers

Arun
5 years ago
Dear student
 
 
P=FV
P=mav
P=mv . dv/ds . v
P/m.ds =v^2.dv
Integrate both sides under limits
V from v1 to v2
ds from 0 to S
We get
(v^3/3)v1tov2=P/m.s
1/3(v2^3-v1^3)=P/m.S
M/3P(v2^3-v1^3)=S
 
Regards
Arun